OpenStudy (anonymous):

Parts for a car repair cost $175. The mechanic charges $34 per hour. If you receive an estimate for at LEAST $226 and at MOST $294 for fixing the car, what is the time interval that the mechanic will be working on the job?

OpenStudy (anonymous):

is it supposed to be set up like this?

OpenStudy (anonymous):

\[75+34x \le226 \]

OpenStudy (anonymous):

\[175+34x \le 294\]

OpenStudy (anonymous):

@completeidiot

OpenStudy (anonymous):

175+34 t where t is the number of hours he worked must be at least 224 and at most 294 your 2nd equation makes sense, your first equation, the sign is the wrong way

OpenStudy (anonymous):

putting it all together in a single equation \[224\le 175+34t\le294\]

OpenStudy (anonymous):

okay, but does it look correct?the way i did it?

OpenStudy (anonymous):

read the first comment i wrote

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

your method is correct, but its quicker to solve it the way i have it written

OpenStudy (anonymous):

\[1.5lex \ge3.5\]so my final answer should look like this

OpenStudy (anonymous):

didnt do the problem, assumeing your math i correct, then the answer would be between time 1.5 h and 3.5 h

OpenStudy (anonymous):

\[1.5\le x \ge 3.5\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

note the sign, what you have written is x greater than 1.5 x is greater than 3.5 you want a range,

OpenStudy (anonymous):

what you want is x is greater than 1.5 and x is less than 3.5 so 1.5< x <3.5

OpenStudy (anonymous):

oh okay i got . Thank you